OpenStudy (quickstudent):

How do you factor this polynomial: -16t^2 + 28t + 8 = 0?

OpenStudy (acxbox22):

simplify by dividing by 4 -4t^2+7t+2=0

OpenStudy (acxbox22):

now we just factor (4t +1)(-t + 2 )=0

OpenStudy (bradely):

-16t^2 + 28t + 8 = 0 Divide by -4 4t^2-7t-2=0 4t^2-8t+t-2=0 4t(t-2)+1(t-2)=0 (t-2)(4t+1)=0 t=2,t=-1/4 Source: http://www.mathskey.com/question2answer/